(3)(a) Suppose that the market for a particular type of product purchased is perfectly competitive, what questions would you ask to determine the validity of the claim?

    (b) Explain, with appropriate graphical illustration(s), why firms in a perfectly competitive market would earn zero economic profits in the long run.

 (4) Consider a market with the following demand and supply curves:

                                                 Q (p) = 20 – 2P               

                                                 Q (p) = - 10 + 3P

(a)Find the equilibrium price and the equilibrium quantity transacted in this market.



1
Expert's answer
2021-12-14T11:18:03-0500

A. The number of sellers in the market.

B.

Profits and losses are eliminated in the long run because an infinite number of enterprises produce infinitely divisible, homogeneous items. As a result, in the long run, the model of perfect competition assumes away all plausible drivers of profits.

C.

"Q d=Qs\\ = 20 – 2P =- 10 + 3P\\20+10=3p+2p\\30=5p\\p=6\\Q=20-12=8"